OpenStudy (anonymous):

3 1/5 divided by 1 1/3.

OpenStudy (anonymous):

\[3\frac{1}{5}=\frac{3*5+1}{5}=\frac{16}{5}\] \[1\frac{1}{3}=\frac{1*3+1}{3}=\frac{4}{3}\] \[\frac{\frac{16}{5}}{\frac{4}{3}}\] \[\frac{16}{5}*\frac{3}{4}\] \[\frac{4}{5}*\frac{3}{1}\] \[\frac{12}{5}\]

OpenStudy (anonymous):

\[\frac{3+\frac{1}{5}}{1+\frac{1}{13}}\text{=}\frac{\frac{16}{5}}{\frac{14}{13}}\text{=}\frac{16}{5}\left(\frac{13}{14}\right)\text{=}\frac{104}{35}\]

OpenStudy (anonymous):

@robtobey from where you got that 1/13

OpenStudy (anonymous):

Thank you for catching my blunder. Your answer of 12/5 is the correct answer.